the ratio height: length: depth of this cuboid is 2:3:4
the total surface area is 1300cm^2.
find the volume of the cuboid.

Answer by phillywily(18)

H : L : D
2 : 3 : 4
2y : 3y : 4y
1300 = 2(2y)(3y) + 2(2y)(4y) + 2(3y)(4y)
1300 = 12y² + 16y² + 24y²
1300 = y² (12+15+24)
1300/y² = 52
y² = 1300/52
y = √25
y = 5
Volume
= 10x15x20
= 3000cm²